OpenStudy (safari321):

hat is square root 300 minus square root 3

OpenStudy (anonymous):

\[\sqrt{300}= \sqrt{3*10*10} = 10\sqrt{3}\]

OpenStudy (anonymous):

so \[\sqrt{300} - \sqrt{3} = 10\sqrt{3} - \sqrt{3} = 9\sqrt{3}\]

OpenStudy (safari321):

is that the final answer

OpenStudy (anonymous):

yes it is.....

OpenStudy (anonymous):

unless u want it in decimal.........

OpenStudy (safari321):

no

OpenStudy (safari321):

that's it ,thanks

OpenStudy (safari321):

What is \[\sqrt{108}\] + \[2\sqrt{3}\] - \[\sqrt{75}\]

OpenStudy (anonymous):

root(108) = root(2*2*3*3*3) = 2*3root(3) = 6 root(3) root(75) = root(3*5*5) = 5 root(3)

OpenStudy (safari321):

thank you

OpenStudy (anonymous):

so \[\sqrt{108} + 2\sqrt{3} - \sqrt{75} = 6\sqrt{3} + 2\sqrt{3} - 5\sqrt{3} = 3\sqrt{3}\]

OpenStudy (safari321):

one more question. \[\sqrt{125}-\sqrt{245}-\sqrt{45}\]